Decide whether the pair of lines is parallel, perpendicular, or neither. 4x + 2y = 9 2x - 4y = 8
Answer #1

© 471+27=9 an-ay=8 :7 from @ zy 9-42 > y = -4/22+9/2 =-2019/2 and from 2 ay = 25.-8. = y ₂ - sta = %=2 of two lines BWA In in